OpenStudy (anonymous):

Multiply -7/1 • -2/9. Your answer should be reduced to its simplest form, and should not be an improper fraction. Answer 5/9 -5/9 1 and 5/9 -1 and 5/9

hero (hero):

@irmadi33, have you multiplied the given fractions yet?

OpenStudy (anonymous):

yeah i got 7/5 v.v

hero (hero):

\(-\left(\dfrac{7}{1}\right) \dot\ -\left(\dfrac{2}{9}\right)\) right?

OpenStudy (anonymous):

yes

hero (hero):

In general \(\dfrac{a}{b} \times \dfrac{c}{d} = \dfrac{ab}{cd}\) Therefore \(-\dfrac{7}{1} \times -\dfrac{2}{9} = \dfrac{(-7)(-2)}{(1)(9)}\)

hero (hero):

So what is the result of that multiplication?

OpenStudy (anonymous):

14/9?

hero (hero):

Correct. And now we must convert the improper fraction to a mixed fraction. Notice that \[\frac{14}{9} = \frac{9 + 5}{9}\]

hero (hero):

Do you notice anything we can do as a result of writing the fraction in that form?